# Numerical valuation of European options under two-asset infinite-activity exponential Lévy models

### v1

@@ Abstract
We propose a numerical method for the valuation of European-style options under two-asset infinite-activity exponential Lévy models. Our method extends the effective approach developed by Wang et al. (2007) for the 1-dimensional case to the 2-dimensional setting and is applicable for general Lévy measures under mild assumptions. A tailored discretization of the non-local integral term is developed, which can be efficiently evaluated by means of the fast Fourier transform. For the temporal discretization, the semi-Lagrangian $\theta$-method is employed in a convenient splitting fashion, where the diffusion term is treated implicitly and the integral term is handled explicitly by a fixed-point iteration. Numerical experiments for put-on-the-average options under Normal Tempered Stable dynamics reveal favourable second-order convergence of our method whenever the exponential Lévy process has finite-variation.
@@ 1 Introduction
The accurate valuation of derivative securities in modern financial markets requires modeling techniques capable of capturing empirical irregularities in asset price dynamics. Classical models based on Brownian motion, such as the Black–Scholes model, rely on continuous-path diffusion and fail to reflect important stylized facts, such as heavy tails and skewness in log-returns. This has motivated the use of Lévy processes in the last decades, which naturally offers a richer class of models for asset dynamics. Among various Lévy models, the Normal Inverse Gaussian (NIG) process has emerged as a parsimonious and effective choice to capture such characteristics. Among others, Rydberg (1997) shows how the NIG model provides a significantly better statistical fit to equity return data compared to classical Gaussian-based models. Lévy models allow for a more realistic representation of market risk and are therefore natural candidates for use in option pricing models.
@@
In this paper, we propose a numerical method for pricing European-style financial derivatives written on two underlying assets, whose dynamics are driven by a 2-dimensional Lévy process, with particular focus on infinite activity processes. Financial pricing under jump-diffusion models can be approached through various methodologies, such as Monte Carlo simulation, Fourier-based methods, and partial integro-differential equations (PIDEs). Monte Carlo methods are flexible and easy to implement, but they suffer from slow convergence. Fourier-based methods, such as in Jackson et al. (2008) and Ruijter & Oosterlee (2012), can be applied when the characteristic exponent of the process is known, and they can achieve exponential convergence. Numerical methods for PIDEs, such as in Cont & Voltchkova (2005), d’Halluin et al. (2005), Wang et al. (2007), Clift & Forsyth (2008), Salmi et al. (2014), Kaushansky et al. (2018), Boen & in ’t Hout (2021) and in ’t Hout & Lamotte (2023), can instead be applied when the Lévy measure is known, and do not require knowledge of the characteristic exponent. They are applicable to a wide variety of financial derivatives.
@@
The numerical method derived in this paper focuses on the case where the underlying 2-dimensional Lévy process exhibits infinite-activity, meaning that an infinite number of jumps occur over any finite time horizon. In this setting, particular care must be taken in the discretization of the non-local 2-dimensional integral term in the PIDE near the origin, where the Lévy measure becomes singular. Indeed, classical quadrature formulae fail to yield the desired second-order convergence.
@@
The main contribution of this paper is an extension of the effective numerical solution approach of Wang et al. (2007) from the 1-dimensional to the 2-dimensional setting. Here a key idea, originally introduced in Asmussen & Rosiński (2001) and Cont & Voltchkova (2005), is to replace the small jumps with an artificial diffusion term. This substitution enables the development of a tailored quadrature scheme, which restores the desired order of convergence of the entire numerical scheme. For the efficient evaluation of the discretized integral operator, a fast Fourier transform (FFT) algorithm is constructed. For the temporal discretization, the semi-Lagrangian $\theta$-method is considered. Here operator splitting is applied, where the diffusion term is treated implicitly and the integral term is handled explicitly by a fixed-point iteration. For the large linear system in each time step, the BiCGSTAB iterative solver is used.
@@
An outline of this paper is as follows. In Section 2, we introduce the market model and the PIDE for the derivative pricing. In Section 3 the proposed numerical scheme is derived. Numerical experiments are discussed in Section 4. The final Section 5 contains our conclusions.
@@ 2.1 Market model
Let $\left(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t\in\left[0,T\right]},\mathbb{P}\right)$ be a filtered probability space, for some given $T>0$. We consider an arbitrage free market characterized by a constant (instantaneous) risk-free interest rate $r$ and an equivalent martingale measure $\mathbb{Q}\sim\mathbb{P}$. We assume that there exist two risky assets, whose prices are modeled by the 2-dimensional stochastic process $X=\left(X^{\left(1\right)},X^{\left(2\right)}\right)$ that solves the following stochastic differential equation
@@
The functions $\mu:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2}$, $\Sigma:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2\times 2}$ and $\gamma:\mathbb{R}^{2}\times\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2}$ are called drift, diffusion and jump function (or term) respectively, where $\mathbb{R}_{\geq 0}^{2}=\left\{x\in\mathbb{R}^{2}:x^{\left(i\right)}\geq 0\text{ for }i=1,2\right\}$. In this paper, we consider the case of the well-known exponential Lévy process, i.e. where the coordinates of the coefficient functions are defined as follows
@@
Let $\left\|\cdot\right\|$ be any given norm on $\mathbb{R}^{2}$. In this paper we assume that $\ell$ is absolutely continuous, has finite variance, i.e.
$\int_{\mathbb{R}}z^{2}\ell\left(dz\right)<\infty$
and there exist constants $A_{\ell}<2$, $B_{\ell}>2$ such that for any given $h>0$ there is $C_{\ell}\left(h\right)>0$ such that
$\begin{cases}\ell\left(z\right)\leq C_{\ell}\left(h\right)\left\|z\right\|^{-A_{\ell}-2}&\text{for any }z\text{ such that }\left\|z\right\|\in\left(0,h\right]\\ \ell\left(z\right)=O\left(e^{-B_{\ell}\left\|z\right\|}\right)&\text{as }\left\|z\right\|\rightarrow\infty.\end{cases}$ (2.5)
The number $A_{\ell}$ governs the activity and variation of the process: $X$ is of finite-activity if $A_{\ell}<0$, since $\int_{\mathbb{R}^{2}}\ell\left(dz\right)<\infty$; it is of finite-variation if $A_{\ell}<1$, since $\int_{\left\|z\right\|<\epsilon}\left\|z\right\|\ell\left(dz\right)<\infty$ for any $\epsilon>0$. The number $B_{\ell}$ characterizes the exponential decay of $\ell$ at infinity. Since the process $X$ has finite moments of all orders up to $k\in\mathbb{N}$ if and only if $\int_{\left\|z\right\|>\epsilon}e^{k\left\|z\right\|}\ell\left(dz\right)<\infty$ for any $\epsilon>0$, then $k<B_{\ell}$ provides a necessary condition of it. Following Applebaum (2004, Chapter 6), the stronger condition $B_{\ell}\geq 2$ is necessary to guarantee the existence of a unique solution with finite variance to the stochastic differential equation (2.1). Most of the common Lévy processes in finance satisfy the conditions (2.5), such as the Kou, Carr-Geman-Madan-Yor (CGMY), Variance Gamma (VG) and Normal Inverse Gaussian (NIG) models.
@@ 3.1 Integral discretization
When the Lévy measure is singular, it is not possible to apply classical quadrature formulae, such as the midpoint or the trapezoidal rule. In fact, in this case the error will blow up as the number of quadrature points increases. To address this problem, we develop in this subsection a different quadrature formula.
@@
where $\left\|z\right\|_{\infty}=\max_{j=1,2}\left|z^{\left(j\right)}\right|$ and $0<z_{\max}^{\mathbf{I}}<z_{\max}^{\mathbf{II}}<z_{\max}^{\mathbf{III}}$ are given numbers, which will be specified in Section 4. The above three sets represent a partition of $R_{z}=\left\{z\in\mathbb{R}^{2}:\left\|z\right\|_{\infty}\leq z_{\max}^{\mathbf{III}}\right\}$, which is a square centered at the origin, as shown in Figure 1.
@@
a high level of accuracy is achieved, despite the integrand being singular at the origin. Clearly, the quadrature weights used in $R_{z}^{\mathbf{II}}$ are constructed as integrals of the Lévy measure, which turns out to be beneficial for the convergence behaviour (as $N_{z}\rightarrow\infty$). Analogously to the entries of the matrix $\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)$ in (3.3), the integrals $\int_{R_{l_{1}l_{2}}}\left\|z\right\|^{2}\ell\left(dz\right)$ can be precomputed using a common numerical integrator. Regarding $R_{z}^{\mathbf{III}}$, the coefficients are obtained by applying the classical midpoint rule, see for example Quarteroni et al. (2007). Finally, note that the weights $\omega_{l_{1}l_{2}}$ are null over $R_{z}^{\mathbf{I}}$, as the first integral in (3.2) has already been approximated through (3.3).
Using (3.3) and substituting (2.7) into (3.4), we can define an approximating operator $\mathcal{A}_{\omega}$ and a number $r_{\omega}$ such that
@@
$\mathcal{A}_{\omega}u\left(x,t\right)$ $=\mu_{\omega}\left(x\right)^{\top}u_{x}\left(x,t\right)+\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ\Sigma_{\omega}\Sigma_{\omega}^{\top}\left(x\right)\right)\mathbf{1}+\left(\mathcal{B}_{\omega}u\right)\left(x,t\right)$ (3.6)
$r_{\omega}$ $=r+\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}$
@@
$\mu_{\omega}^{\left(i\right)}\left(x\right)$ $=x^{\left(i\right)}\kappa_{\omega}^{\left(i\right)}$
$\kappa_{\omega}^{\left(i\right)}$ $=r-\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}\left(e^{z_{l_{1}l_{2}}^{\left(i\right)}}-1\right)$
$\left(\Sigma_{\omega}\Sigma_{\omega}^{\top}\right)^{\left(i,j\right)}\left(x\right)$ $=x^{\left(i\right)}x^{\left(j\right)}\left(\sigma_{\omega}\sigma_{\omega}^{\top}\right)^{\left(i,j\right)}$
$\sigma_{\omega}\sigma_{\omega}^{\top}$ $=\sigma\sigma^{\top}+\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)$
@@ 3.2 Spatial discretization
Let $R_{x}=\left[0,x_{\max}\right]\times\left[0,x_{\max}\right]$ be the truncated $x$-domain over which the solution to (3.8) is approximated and $N_{x}\in\mathbb{N}$ be a given number of spatial grid points. Here, $x_{\max}$ is heuristically chosen in such a way that the localization error is negligible. We construct a spatial grid $\mathbf{x}$ in $R_{x}$ by applying, in each dimension, a strictly increasing and smooth transformation $\varphi$ to an artificial uniform grid. Let
$x_{m}=\varphi\left(\varphi^{-1}\left(0\right)+\frac{\varphi^{-1}\left(x_{\max}\right)-\varphi^{-1}\left(0\right)}{N_{x}}m\right)\qquad\left(m=0,1,\ldots,N_{x}\right)$
with $h_{x,m}=x_{m}-x_{m-1}$. The elements of $\mathbf{x}$ are defined by

### v2

@@ Abstract
We propose a numerical method for the valuation of European-style options under two-asset infinite-activity exponential Lévy models. Our method extends the effective approach developed by Wang et al. (2007) for the 1-dimensional case to the 2-dimensional setting and is applicable for general Lévy measures under mild assumptions. A tailored discretization of the non-local integral term is developed, which can be efficiently evaluated by means of the fast Fourier transform. For the temporal discretization, the semi-Lagrangian $\theta$-method is employed in a convenient splitting fashion, where the diffusion term is treated implicitly and the integral term is handled explicitly by a fixed-point iteration. Numerical experiments for put-on-the-average options under Normal Tempered Stable dynamics reveal a favourable convergence behaviour of our method whenever the exponential Lévy process has finite-variation. In addition, a relevant theoretical convergence result for the discretization of the integral term is proved.
@@ 1 Introduction
The accurate valuation of derivative securities in modern financial markets requires modeling techniques capable of capturing empirical irregularities in asset price dynamics. Classical models based on Brownian motion, such as the Black–Scholes model, rely on continuous-path diffusion and fail to reflect important stylized facts, such as heavy tails and skewness in log-returns. This has motivated the use of Lévy processes in the last decades, which naturally offer a richer class of models for asset dynamics. Among various Lévy models, the Normal Inverse Gaussian (NIG) process has emerged as a parsimonious and effective choice to capture such characteristics. Among others, Rydberg (1997) shows how the NIG model provides a significantly better statistical fit to equity return data compared to classical Gaussian-based models. Lévy models allow for a more realistic representation of market risk and are therefore natural candidates for use in option pricing models.
@@
In this paper, we propose a numerical method for pricing European-style financial derivatives written on two underlying assets, whose dynamics are driven by a 2-dimensional Lévy process, with particular focus on infinite-activity processes. Financial pricing under jump-diffusion models can be approached through various methodologies, such as Monte Carlo simulation, Fourier-based methods, and partial integro-differential equations (PIDEs). Monte Carlo methods are flexible and easy to implement, but they suffer from slow convergence. Fourier-based methods, such as in Jackson et al. (2008) and Ruijter & Oosterlee (2012), can be applied when the characteristic exponent of the process is known, and they can achieve exponential convergence. Numerical methods for PIDEs, such as in Cont & Voltchkova (2005), d’Halluin et al. (2005), Wang et al. (2007), Clift & Forsyth (2008), Salmi et al. (2014), Kaushansky et al. (2018), Boen & in ’t Hout (2021) and in ’t Hout & Lamotte (2023), can instead be applied when the Lévy measure is known, and do not require knowledge of the characteristic exponent. They are applicable to a wide variety of financial derivatives.
@@
The numerical method derived in this paper focuses on the case where the underlying 2-dimensional Lévy process exhibits infinite-activity, meaning that an infinite number of jumps occur over any finite time horizon. In this setting, particular care must be taken in the discretization of the non-local 2-dimensional integral term in the PIDE near the origin, where the Lévy measure becomes singular.
@@
The main contribution of this paper is an extension of the effective numerical solution approach of Wang et al. (2007) from the 1-dimensional to the 2-dimensional setting. Here, a key idea, originally introduced in Asmussen & Rosiński (2001) and Cont & Voltchkova (2005), is to replace the small jumps with an artificial diffusion term. This substitution enables the development of a tailored quadrature scheme. For the efficient evaluation of the discretized integral operator, a fast Fourier transform (FFT) algorithm is constructed. For the temporal discretization, the semi-Lagrangian $\theta$-method is considered. Here, operator splitting is applied, where the diffusion term is treated implicitly and the integral term is handled explicitly by a fixed-point iteration. For the large linear system in each time step, the BiCGSTAB iterative solver is used.
To assess the performance of the proposed numerical method, we conduct numerical experiments for put-on-the-average options under Normal Tempered Stable dynamics. These experiments reveal a favourable convergence behaviour whenever the underlying exponential Lévy process has finite-variation. A relevant theoretical result is proved on the order of convergence for the discretization of the integral term.
@@
An outline of this paper is as follows. In Section 2, we introduce the market model and the PIDE for the derivative pricing. In Section 3 the proposed numerical scheme is derived and a theorem on the approximation error for the integral term is presented. Numerical experiments are discussed in Section 4. The final Section 5 contains our conclusions.
@@ 2.1 Market model
Let $\left(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t\in\left[0,T\right]},\mathbb{P}\right)$ be a filtered probability space, for some given $T>0$. We consider an arbitrage free market characterized by a constant (instantaneous) risk-free interest rate $r$ and an equivalent martingale measure $\mathbb{Q}\sim\mathbb{P}$. We assume that there exist two risky assets whose prices are modeled by the 2-dimensional stochastic process $X=\left(X^{\left(1\right)},X^{\left(2\right)}\right)$ that solves the following stochastic differential equation
@@
The functions $\mu:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2}$, $\Sigma:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2\times 2}$ and $\gamma:\mathbb{R}^{2}\times\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2}$ are called drift, diffusion, and jump function (or term) respectively, where $\mathbb{R}_{\geq 0}^{2}=\left\{x\in\mathbb{R}^{2}:x^{\left(i\right)}\geq 0\text{ for }i=1,2\right\}$. In this paper, we consider the case of the well-known exponential Lévy process, i.e. where the coordinates of the coefficient functions are defined for $i,j=1,2$ as follows
@@
Let $\left\|\cdot\right\|$ be any given norm on $\mathbb{R}^{2}$. In this paper we assume that $\ell$ is absolutely continuous, the corresponding Lévy process has finite variance, i.e.
$\int_{\mathbb{R}^{2}_{*}}\left\|z\right\|^{2}\ell\left(dz\right)<\infty,$ (2.5)
and there exist constants $A_{\ell}<2$, $B_{\ell}>2$ and $C_{\ell}\left(h\right)$, $C_{\ell}^{\prime}\left(h\right)>0$, for any $h>0$, such that
[…]
where $\ell_{z}^{\left(j\right)}$ denotes the $j$-th partial derivative of $\ell$ with respect to $z$, with $j=1,2$. The number $A_{\ell}$ governs the activity and variation of the process: $X$ is of finite-activity if $A_{\ell}<0$, since $\int_{\mathbb{R}^{2}_{*}}\ell\left(dz\right)<\infty$; it is of finite-variation if $A_{\ell}<1$, since $\int_{0<\left\|z\right\|<\epsilon}\left\|z\right\|\ell\left(dz\right)<\infty$ for any $\epsilon>0$. The number $B_{\ell}$ characterizes the exponential decay of $\ell$ at infinity. Since the process $X$ has finite moments of all orders up to $k\in\mathbb{N}$ if and only if $\int_{\left\|z\right\|>\epsilon}e^{k\left\|z\right\|}\ell\left(dz\right)<\infty$ for any $\epsilon>0$, then $k<B_{\ell}$ provides a necessary condition of it. Following Applebaum (2004, Chapter 6), the stronger condition $B_{\ell}\geq 2$ is necessary to guarantee the existence of a unique solution with finite variance to the stochastic differential equation (2.1). Most of the common Lévy processes in finance satisfy the conditions (2.6), such as the Kou, Carr–Geman–Madan–Yor (CGMY), Variance Gamma (VG) and Normal Inverse Gaussian (NIG) models.
@@ 3.1 Integral discretization
When the Lévy measure is singular, classical quadrature formulae such as the midpoint rule or the trapezoidal rule fail. In fact, in this case the error will blow up as the number of quadrature points increases. To address this problem, we develop in this subsection a different quadrature formula.
@@
where $\left\|z\right\|_{\infty}=\max_{j=1,2}\left|z^{\left(j\right)}\right|$ and $0<z_{\max}^{\mathbf{I}}<z_{\max}^{\mathbf{II}}<z_{\max}^{\mathbf{III}}$ are given numbers, which will be specified below. The above three sets represent a partition of $R_{z}=\left\{z\in\mathbb{R}^{2}:\left\|z\right\|_{\infty}\leq z_{\max}^{\mathbf{III}}\right\}$, which is a square centered at the origin, as shown in Figure 1.
@@
a satisfactory level of accuracy is achieved, despite the integrand being singular at the origin. Clearly, the quadrature weights used in $R_{z}^{\mathbf{II}}$ are constructed as integrals of the Lévy measure, which turns out to be beneficial for the convergence behaviour (as $N_{z}\rightarrow\infty$). Analogously to the entries of the matrix $\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)$ in (3.3), the integrals $\int_{R_{l_{1}l_{2}}}\left\|z\right\|^{2}\ell\left(dz\right)$ can be precomputed using a common numerical integrator. Regarding $R_{z}^{\mathbf{III}}$, the coefficients are obtained by applying the classical midpoint rule, see for example Quarteroni et al. (2007). Finally, note that the weights $\omega_{l_{1}l_{2}}$ are null over $R_{z}^{\mathbf{I}}$, as the first integral in (3.2) has already been approximated through (3.3).
For the above approximation of the integral term, we have the following convergence result.
Proposition 3.1.
Let $u$ (resp. $\ell$) be sufficiently smooth with respect to $x$ (resp. $z$). Let $z_{\max}^{\mathbf{I}}$ be directly proportional to $h_{z}$ and let $z_{\max}^{\mathbf{II}}$, $z_{\max}^{\mathbf{III}}$ be fixed (independent of $h_{z}$). Let $\left\|\cdot\right\|$ be the Euclidean norm. Assume $h_{z}$ is such that $z_{\max}^{\mathbf{II}}/h_{z}$ is an integer. Then, for any given $\epsilon>0$ and $x_{\max}>0$, the quadrature error
$E\left(x,t\right)$ $=\int_{\left\|z\right\|_{\infty}<z_{\max}^{\mathbf{III}}}f\left(z,x,t\right)\ell\left(dz\right)$ (3.6)
$\quad-\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ I_{x}\left(\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)\right)I_{x}\right)\mathbf{1}-\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}f\left(z_{l_{1}l_{2}},x,t\right)$
satisfies
$E\left(x,t\right)=\begin{cases}O\left(h_{z}^{2-A_{\ell}}\right)&\text{if }0<A_{\ell}<2,\\ O\left(h_{z}^{2-\epsilon}\right)&\text{if }A_{\ell}=0,\end{cases}$
uniformly in $\left(x,t\right)\in\left[0,x_{\max}\right]^{2}\times\left[0,T\right]$.
Proof.
See Appendix B. ∎
Proposition 3.1 corresponds to a result derived by Wang et al. (2007) for the 1-dimensional setting. Clearly, in our 2-dimensional setting, essentially second-order convergence is obtained for $A_{\ell}=0$ and the order of convergence equals $2-A_{\ell}$ whenever $0<A_{\ell}<2$. This forms a positive result. We note, however, that it is less advantageous compared to the 1-dimensional setting, as in this case Wang et al. (2007) proved second-order convergence whenever $0\leq A_{\ell}<1$, essentially second-order convergence if $A_{\ell}=1$, and an order of convergence equal to $3-A_{\ell}$ whenever $1<A_{\ell}<2$. The less advantageous convergence result in the 2-dimensional setting is attributed to a lack of smoothness w.r.t. $z$ of the scaled function $f\left(z,x,t\right)\left\|z\right\|^{-2}$ in the neighbourhood of $z=0$, see the proof in Appendix B.
Using (3.3) and substituting (2.8) into (3.4), we define an approximating operator $\mathcal{A}_{\omega}$ and a number $r_{\omega}$ such that
@@
$\mathcal{A}_{\omega}u\left(x,t\right)$ $=\mu_{\omega}\left(x\right)^{\top}u_{x}\left(x,t\right)+\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ\Sigma_{\omega}\Sigma_{\omega}^{\top}\left(x\right)\right)\mathbf{1}+\left(\mathcal{B}_{\omega}u\right)\left(x,t\right),$ (3.7)
$r_{\omega}$ $=r+\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}},$
@@
$\mu_{\omega}^{\left(i\right)}\left(x\right)$ $=x^{\left(i\right)}\kappa_{\omega}^{\left(i\right)},$
$\kappa_{\omega}^{\left(i\right)}$ $=r-\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}\left(e^{z_{l_{1}l_{2}}^{\left(i\right)}}-1\right),$
$\left(\Sigma_{\omega}\Sigma_{\omega}^{\top}\right)^{\left(i,j\right)}\left(x\right)$ $=x^{\left(i\right)}x^{\left(j\right)}\left(\sigma_{\omega}\sigma_{\omega}^{\top}\right)^{\left(i,j\right)},$
$\sigma_{\omega}\sigma_{\omega}^{\top}$ $=\sigma\sigma^{\top}+\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right),$
@@ 3.2 Spatial discretization
Let $R_{x}=\left[0,x_{\max}\right]\times\left[0,x_{\max}\right]$ be the truncated $x$-domain over which the solution to (3.9) is approximated and $N_{x}\in\mathbb{N}$ be a given number of spatial grid points. Here, $x_{\max}$ is chosen heuristically in such a way that the localization error is negligible. We construct a spatial grid $\mathbf{x}$ in $R_{x}$ by applying, in each dimension, a strictly increasing and smooth transformation $\varphi$ to an artificial uniform grid. Let
$x_{m}=\varphi\left(\varphi^{-1}\left(0\right)+\frac{\varphi^{-1}\left(x_{\max}\right)-\varphi^{-1}\left(0\right)}{N_{x}}m\right)\qquad\left(m=0,1,\ldots,N_{x}\right).$
The elements of $\mathbf{x}$ are defined by